# TabSTAR: A Foundation Tabular Model With Semantically Target-Aware Representations

### v1

@@ 5 Results
We evaluate each model using AUROC (classification) and $R^{2}$ (regression) as metrics. Following [37], we normalize scores per dataset split to the $[0,1]$ range, using the best and worst model performance as anchors. [^15] For a single run, the best model gets 1, the worst gets 0 and the rest are linearly scaled accordingly. The normalized scores are averaged across all runs, with 95% CIs. Performance for all models on both conditions are shown in Figure 2 (classification) and Figure 3 (regression). Besides their example limit, TabPFN-v2 cannot process 4 datasets (more than 10 target classes or more than 500,000 cells) and CARTE cannot handle 15 (a bug in their PowerTransformer implementation). Reported averages for these models are computed only over the datasets where evaluation is feasible. Appendix F expands on dataset level performance, head-to-head comparisons and running times.
@@
In classification problems, TabSTAR consistently achieves SOTA performance. This is evident both when restricting the dataset size to 10,000 examples and when using larger datasets in the unlimited condition. For the 10K condition, TabSTAR achieves a 0.809 score, performing better than TabPFN-v2 (0.783) and significantly better than GBDTs (0.756 CatBoost-Tuned, 0.744 XGB-Tuned). When analyzing head-to-head comparisons (Appendix F.2), TabSTAR outperforms TabPFN-v2 (7/11 datasets), XGB-Tuned (10/14) and CatBoost-Tuned (11/14). For the Unlimited condition, TabSTAR-Unlimit achieves a 0.874 score, significantly above the second-best CatBoost-Tuned-Unlimit with 0.734. Importantly, all Unlimit variants surpass the 10K ones, emphasizing the importance of scaling.
@@
Although regression is not our main focus, TabSTAR achieves competitive results in the 10K condition, but clearly does not set the SOTA. Surprisingly, while TabPFN-v2 is superior, it significantly underperforms compared to GBDTs which dominate this category. This emphasizes the need for better modeling of textual tabular learning, especially since TabPFN-v2 has shown remarkable performance in non-textual tabular datasets, and CARTE set the SOTA for small datasets. When analyzing the Unlimited variants, TabSTAR scales well, and surpasses other TFMs which cannot scale, but the gap from GBDTs remains significant. §7 discusses this limitation and suggests promising directions for future generations of TabSTAR to achieve SOTA in regression as well.
@@ 6 Analysis
To answer these questions, we pretrain several variants of TabSTAR for each analysis, limiting ourselves to a subset of the tabular datasets used for the main experiment (see §4). Specifically, each variant is pretrained over 256 datasets [^16] Except for variants of Q2, which analyze the effect of number of datasets on pretraining. including 30 datasets from our benchmark, and evaluated over the remaining 20 datasets (12 regression, 8 classification). This reduced setup allows leveraging transfer learning and exploiting our corpus, without the burden of training multiple folds per variant. Results are reported with the same normalized metric used in §5, scaling the performance to the $[0,1]$ range. Appendix G.1 lists the 20 datasets used for evaluation along with per-dataset results.
@@
Q1: The Role of the Encoder Unfreezing
To investigate whether unfreezing layers of the textual encoder impacts performance, we conduct experiments where we unfreeze varying numbers of the 12 encoder layers during both TabSTAR’s pretraining and finetuning stages. [^17] For each variant, the number of unfrozen layers remains the same in both pretraining and finetuning. Figure 4 shows the validation loss during TabSTAR pretraining (left) and the normalized score on the downstream tasks (right) as a function of the number of unfrozen encoder layers. Notably, unfreezing even a single encoder layer significantly outperforms using static embeddings. Further substantial improvements are observed as more layers are tuned, with the best results achieved when unfreezing 6 layers. While unfreezing 9 layers shows lower performance, it is plausible that adding more datasets to the pretraining phase will affect this finding. See Appendix G.2 for more details.
Pretraining Datasets 0 16 64 256
Classification 0.352 ± 0.086 0.450 ± 0.084 0.558 ± 0.086 0.786 ± 0.076
Regression 0.338 ± 0.073 0.395 ± 0.068 0.642 ± 0.066 0.811 ± 0.055
@@
To evaluate the impact of pretraining on TabSTAR’s downstream performance, we compare a pretrained version of TabSTAR with a version that was finetuned from scratch. [^18] Since LoRA underperforms on random weights, we finetune the entire non-pretrained model. In line with previous work [86, 82], the pretrained model performs significantly better, highlighting the critical role of transfer learning for TabSTAR’s success. To further investigate the effect of the number of pretraining datasets on downstream task performance, we train two additional versions: one pretrained on 16 datasets and another on 64 datasets. As shown in Table 3, increasing the number of pretraining datasets consistently improved performance in both classification and regression tasks. Notably, the substantial gain in regression tasks suggests that TabSTAR’s downstream performance on §5 could improve with more pretraining data (see Appendix G.3).
@@
A key challenge in integrating language models with numerical data is determining how to best represent numerical values within a linguistic framework. While some semantic tabular methods omit numerical features from the verbalization [79, 82], TP-BERTa [80] introduced Relative Magnitude Tokenization [80], which encode numerical information through non-semantic special bin tokens. In constrast, TabSTAR injects semantic numerical information into the verbalization of numerical features, as illustrated in Table 2. To quantify the effect of our novel verbalization, we explore two thinner variants: (1) Name + Bin, which excludes the quantile information, and (2) Name, which omits numeric information entirely and verbalizes the feature name only. Appendix G.4 shows an illustrative example for each variant and presents the full results. As demonstrated in Table 4, our findings reveal that incorporating numerical information significantly enhances performance, highlighting the importance of balancing numerical precision with a representation format that aligns with the language model’s parametric knowledge.
@@ 7 Discussion and Conclusion
Since scaling laws in data and model size have proven themselves for LLMs [43] and TabSTAR improves with the number of pretraining datasets (see §6), future work should scale TabSTAR across both model and data dimensions. For model scaling, we envision a family of model sizes, common for LLMs [29, 71, 49], that will allow a trade-off between quality and costs. Data scaling might leverage self-supervised learning [59] over large-scale table corpora [17], or realistic synthetic tabular data generators [9], which have proven successful [37, 2]. At scale, it could potentially unlock few-shot learning capabilities and develop automatic feature-engineering skills [36].
Beyond scaling, TabSTAR’s semantic approach has tremendous potential to explicitly include world knowledge, by leveraging LLMs which to date have had limited impact on tabular learning. As a few motivating examples, LLMs could improve TabSTAR’s numerical verbalization binning approach by providing semantically informed thresholds, or provide explicit, contextual world knowledge that could be injected as a strong prior in small data scenarios. While these directions seem like plausible research paths, they come with a risk of data leakage due to the memorization properties of LLMs [8]. Evaluating TFMs fairly while keeping benchmarks uncontaminated would be an important enabler for tabular research. As a step in this direction, we are releasing several TabSTAR variants, each with a different dataset withheld during pretraining, ensuring that for every dataset there is a TabSTAR model that has never seen it. We urge fellow researchers to adopt this approach in their own work.
While TabSTAR sets a new bar in classification, its regression results lag behind GBDTs, which outperform other TFMs as well. This gap could be narrowed through additional scaling, and also by exploring regression-via-classification techniques like [37, 2]. Furthermore, TabSTAR has not been extensively evaluated in few-shot scenarios and in purely numerical datasets. [^19] Partly because of the computational burden of tuning baselines, and the lack of objective leaderboards [73]. In addition, it demands more compute compared to GBDTs, and it may struggle with memory constraints on datasets containing hundreds of features. Despite these limitations, TabSTAR offers a promising pathway toward improving performance on tabular datasets with textual fields, common in industries with high social impact (e.g., healthcare, education), or with significant economic value (e.g., banking, manufacturing). We believe TabSTAR paves the way for a new generation of semantically enriched tabular models, and we welcome the research community’s innovations built on this foundation.
@@
We thank Omri Feldman for brainstorming since the very beginning; Elad Hoffer and Ofir Lindenbaum for consulting and feedback; David Holzmüller and Myung Kim for supporting evaluations; and Noah Hollmann, Léo Grinsztajn, and the Prior Labs team for providing extensive access to TabPFN-v2.
@@ A.1 The Verbalization Module
[…]
@@ A.2 The Annotated TabSTAR
The fused sequence of $e$ tokens is processed by a standard Transformer Encoder with model dimension $d=384$, $L=6$ layers, $6$ attention heads per layer, feed-forward size $4d$, dropout $0.1$, ReLU activation and using a pre-norm configuration. Unlike in language modeling, feature ordering is irrelevant, so no positional encodings are used. The encoder produces contextualized embeddings for every position, and we retain the $t$ target embeddings for the prediction.
@@ A.3 Architecture Experiments
For the fusion block, we consider two simpler alternatives to the attention mechanism, both of then underperforming: (1) Concatenation, by concatenating the semantic and numerical $d$-dimensional vectors into a $2d$-dimensional vector, and projecting them back via an MLP, and (2) Multiplication, by multiplying the semantic representation directly with the numerical value [^22] After rescaling it to be centered around 1 rather than 0, using a learned scaling factor. in a straightforward, parameter-free manner as in [79, 82].
@@ B.2 Finetuning
We finetune downstream tasks using LoRA’s implementation of the peft package [^24] https://github.com/huggingface/peft. We use a rank of $r=32$, set $\alpha=2r=64$ and $dropout=0.1$. We employ the same scheduler as in the pretraining phase, with the only difference being that we set $lr=0.001$, and increase the patience parameter for early stopping to 5. We apply a train-test split of 90%-10% and sample a validation set of 10%. As opposed to the pretraining, all batches are drawn from the same dataset. Therefore, we observe no effect from changing the mini-batch size when keeping the global batch size fixed to 128. We tune 1,597,440 out of TabSTAR’s 47,263,874 parameters (3.4%).
@@ E.2 Baselines Implementation and Hyperparameters
CARTE natively supports textual inputs, and the TabPFN-v2 API client [^28] https://github.com/PriorLabs/tabpfn-client does as well, although its implementation details remain undisclosed. On the other hand, GBDTs do not natively support free-text features, [^29] CatBoost includes a built-in text module, but it underperforms compared to dense text embeddings., so we preprocess these features into fixed-size embeddings using skrub’s TextEncoder, which internally applies a frozen e5-small-v2 encoder to each semantic column. This aligns with the encoder used in TabSTAR, enabling a fair comparison across models. There are, however, two key differences in how TabSTAR handles these embeddings: First, the embeddings are specifically finetuned for the task, contributing significantly to its strong performance as shown in §5 and further analyzed in §G.2. The second detail is that skrub applies dimensionality reduction to 30 dimensions, as proposed by [31]. This compressed representation performs comparably to the full embedding space, while offering improved inference efficiency.
@@ E.2.1 TabPFN-v2
E.2.2 CARTE

### v2

@@ 5 Results
We evaluate each model using AUROC (classification) and $R^{2}$ (regression) as metrics. Following [38], we normalize scores per dataset split to the $[0,1]$ range, using the best and worst model performance as anchors. [^16] For a single run, the best model gets 1, the worst gets 0, and the rest are linearly scaled accordingly. The normalized scores are averaged across all runs, with 95% CIs. Performance for all models on both conditions is shown in Figure 2 (classification) and Figure 3 (regression). For certain datasets, we are unable to execute TabPFN-v2, CARTE, and TabDPT due to model-specific implementation issues or scalability constraints. Reported averages for these models are computed only over the datasets where evaluation is feasible. Appendix F.1 elaborates on technical limitations, Appendix F.2 on dataset-level performance, and Appendix F.3 on head-to-head comparisons.
@@
In classification problems, TabSTAR consistently achieves SOTA performance. This is evident both when restricting the dataset size to 10,000 examples and when using larger datasets in the unlimited condition. For the 10K condition, TabSTAR achieves a 0.83 score, followed by TabM-Tuned (0.73), CatBoost-Tuned (0.69), and TabPFN-v2 (0.67). When analyzing head-to-head comparisons (see Appendix F.3), TabSTAR outperforms TabPFN-v2 (8/11 datasets), TabM-Tuned (10/14), and CatBoost-Tuned (12/14). For the unlimited condition, TabSTAR-Unlimited achieves a 0.84 score, followed by TabM-Tuned-Unlimited (0.79), and much above the rest. Importantly, unlimited variants significantly surpass the 10K ones, emphasizing the importance of models with no scaling restrictions.
@@
Although regression is not our main focus, TabSTAR achieves competitive results in the 10K condition, but clearly does not set the SOTA. Surprisingly, while TabPFN-v2 is superior, it significantly underperforms compared to GBDTs which dominate this category. This emphasizes the need for better modeling of textual tabular learning, especially since TabPFN-v2 has shown remarkable performance in non-textual tabular datasets, and CARTE set the SOTA for small datasets. In the unlimited setting, TabSTAR scales well and surpasses other TFMs which cannot scale, but the gap from GBDTs remains significant. §7 discusses this limitation and suggests promising directions for future generations of TabSTAR to achieve SOTA in regression as well.
Cost Analysis
Table 3 reports the computational requirements of TabSTAR and competing baselines, measured by runtime and peak memory usage, for both training and inference. To highlight the importance of GPU acceleration when text embeddings are part of the preprocessing, we include a variant of CatBoost restricted to CPU only. We exclude TabPFN-v2, whose performance is probably comparable to TabICL, since it is only accessible via an API, and CARTE, whose extensive tuning requirements and prohibitive training costs forced evaluation on less capable hardware.
We find that TabSTAR incurs modest inference costs, comparable to GPU-accelerated GBDTs, since GPU processing of text embeddings dominates runtime. This is evident from the CatBoost-CPU variant, whose inference is nearly ten times slower than TabSTAR. Moreover, TabICL and TabDPT are orders of magnitude slower and consume far more memory, emphasizing the scalability limitations for ICL-based models. During training, TabSTAR is considerably slower than the baselines. However, while fitting a single GBDT is much faster, this advantage quickly diminishes in practice when hyperparameter tuning is needed or GPU acceleration for text embeddings is not available. Appendix F.4 provides dataset-level runtimes and F.5 analyzes costs as a function of the feature count.
Table 3: Cost Analysis. Median training and inference times and peak memory usage on GPU and CPU, aggregated over the 50 datasets of the benchmark with up to 10,000 examples.
Train Inference
Model Time (s) GPU (GB) CPU (GB) Time (s) GPU (GB) CPU (GB)
CatBoost-CPU 360.8 2.2 34.0 2.1
CatBoost 68.5 1.3 1.5 2.0 1.3 1.5
LightGBM 39.1 1.3 1.5 2.0 1.3 1.5
RandomForest 86.1 1.3 1.5 2.5 1.3 1.5
RealMLP 136.4 1.4 1.7 2.5 1.3 1.7
TabDPT 10.4 2.2 1.8 161.4 33.8 6.0
TabICL 42.6 1.2 1.7 13.4 25.4 2.1
TabM 19.1 1.8 4.4 1.5 1.8 4.3
TabSTAR 493.2 4.7 1.8 3.0 1.3 1.8
XGBoost 38.4 1.3 1.5 2.5 1.3 1.5
@@ 6 Analysis
To answer these questions, we pretrain several variants of TabSTAR for each analysis, limiting ourselves to a subset of the tabular datasets used for the main experiment (see §4). Specifically, each variant is pretrained over 256 datasets [^17] Except for variants of Q2, which analyze the effect of the number of datasets on pretraining. including 30 datasets from our benchmark, and evaluated over the remaining 20 datasets (12 regression, 8 classification). This reduced setup allows leveraging transfer learning and exploiting our corpus, without the burden of training multiple folds per variant. Appendix G.1 lists the 20 datasets used for evaluation along with per-dataset results.
Q1: The Role of the Encoder Unfreezing
We examine the effect of unfreezing different numbers of textual encoder layers during pretraining. In finetuning, we keep all base layers frozen and instead apply LoRA adapters to the same layers that were unfrozen in pretraining. [^18] For simplicity, we will refer to them as unfrozen to differentiate them from layers without LoRA. Figure 4 shows the validation loss during TabSTAR pretraining (left) and the normalized score on the downstream tasks (right) as a function of the number of unfrozen encoder layers. Notably, unfreezing even a single encoder layer significantly outperforms using static embeddings. Further substantial improvements are observed as more layers are tuned, with the best results achieved when unfreezing 6 layers. While unfreezing 9 layers shows lower performance, it is plausible that adding more datasets to the pretraining phase will affect this finding. See Appendix G.2 for more details.
@@
To evaluate the impact of pretraining on TabSTAR’s downstream performance, we compare a pretrained version of TabSTAR with a version that was finetuned from scratch. [^19] Since LoRA underperforms on random weights, we finetune the entire non-pretrained model. In line with previous work [90, 86], the pretrained model performs significantly better, highlighting the critical role of transfer learning for TabSTAR’s success. To further investigate the effect of the number of pretraining datasets on downstream task performance, we train two additional versions: one pretrained on 16 datasets and another on 64 datasets.
Figure 5: Average performance on downstream tasks as a function of the number of pretraining datasets (in log scale). We use AUROC for classification (left), and $R^{2}$ for regression (right).
As shown in Figure 5, increasing the number of pretraining datasets consistently improved performance in both classification and regression tasks. Notably, the substantial gain in regression tasks suggests that TabSTAR’s downstream performance on §5 could improve with more pretraining data, potentially reaching SOTA performance with enough scale. See Appendix G.3 for more details.
@@
A key challenge in integrating language models with numerical data is determining how to best represent numerical values within a linguistic framework. While some semantic tabular methods omit numerical features from the verbalization [83, 86], TP-BERTa [84] introduced Relative Magnitude Tokenization [84], which encodes numerical information through non-semantic special bin tokens. In contrast, TabSTAR injects semantic numerical information into the verbalization of numerical features, as illustrated in Table 2. To quantify the effect of our novel verbalization, we explore two thinner variants: (1) Name + Bin, which excludes the quantile information, and (2) Name, which omits numeric information entirely and verbalizes the feature name only. Appendix G.4 shows an illustrative example for each variant and presents the full results. As demonstrated in Table 4, our findings reveal that incorporating numerical information significantly enhances performance, highlighting the importance of balancing numerical precision with a representation format that aligns with the language model’s parametric knowledge.
@@ 7 Discussion and Conclusion
Since scaling laws in data and model size have proven themselves for LLMs [44] and TabSTAR improves with the number of pretraining datasets (see §6), future work should scale TabSTAR across both model and data dimensions. For model scaling, we envision a family of model sizes, common for LLMs [30, 74, 51], that will allow a trade-off between quality and costs. Data scaling might leverage self-supervised learning [62, 82] over large-scale table corpora [18], or realistic synthetic tabular data generators [10], which have proven successful [38, 2]. At scale, it could potentially unlock few-shot learning capabilities and develop automatic feature-engineering skills [37].
Beyond scaling, TabSTAR’s semantic approach has tremendous potential to explicitly include world knowledge by leveraging LLMs, which to date have had a limited impact on tabular learning. As a few motivating examples, LLMs could improve TabSTAR’s numerical verbalization binning approach by providing semantically informed thresholds, or by providing explicit world knowledge that could be injected as a strong prior in small data scenarios. While these directions seem like plausible research paths, they come with a risk of data leakage due to the memorization properties of LLMs [9]. Evaluating TFMs fairly while keeping benchmarks uncontaminated would be an important enabler for tabular research. As a step in this direction, we are releasing several TabSTAR variants, each with a different dataset withheld during pretraining, ensuring that for every dataset there is a TabSTAR model that has never seen it. We urge fellow researchers to adopt this approach in their own work.
While TabSTAR sets a new bar in classification, its regression results lag behind GBDTs, which outperform other TFMs as well. This gap could be narrowed through additional scaling, by exploring regression-via-classification techniques like [38, 2], and by enriching the numerical encoder with distribution-aware statistics per feature as done by [61]. In addition, similar to other TFMs, TabSTAR may encounter memory bottlenecks on datasets with hundreds of features, and its training speed lags behind untuned GBDTs and ICL-based TFMs. Yet TabSTAR achieves GBDT-level efficiency during inference, as opposed to TabPFN-v2 and TabICL. Furthermore, TabSTAR has not been extensively evaluated in few-shot scenarios and in purely numerical datasets.
Despite these limitations, TabSTAR offers a promising pathway toward improving performance on tabular datasets with textual fields, common in industries with high social impact (e.g., healthcare, education), or with significant economic value (e.g., banking, manufacturing). In addition, its architecture lends itself to multimodality and could be extended to tabular datasets that combine numerical, textual, and image features. We believe TabSTAR paves the way for a new generation of semantically enriched, Multimodal TFMs, and invite the research community to advance this vision.
@@
Roi Reichart and Eilam Shapira have been partially supported by a VATAT grant on data science.
We thank Omri Feldman for brainstorming since the very beginning; Elad Hoffer and Ofir Lindenbaum for consulting and feedback; David Holzmüller, Lennart Purucker, Myung Kim, and Gaël Varoquaux for assisting with evaluations and benchmarks; and Frank Hutter, Noah Hollmann, Léo Grinsztajn, and the rest of the Prior Labs team for providing extensive access to TabPFN-v2’s API version.
@@ A.1 The Verbalization Module
Numerical features are represented by both a numerical and a semantic representation. For the numerical representation, given a value $x$, we compute the clipped z-score $z^{\prime}=\operatorname{clip}\bigl((x-\mu)/\sigma,\,-3,\,3\bigr)$ where $\mu,\sigma$ are the training set mean and the standard deviation, and missing values are set to 0. For the semantic representation, we build $B=10$ quantile bins over the training distribution to map the value accordingly. Table 5 shows an illustrative example for the feature Age from our running example in Table 2.
@@ A.2 The Annotated TabSTAR
The fused sequence of $e$ tokens is processed by a standard Transformer Encoder with model dimension $d=384$, $L=6$ layers, $6$ attention heads per layer, feed‐forward size $4d$, dropout $0.1$, ReLU activation and using a pre-norm configuration. Unlike in language modeling, feature ordering is irrelevant, so no positional encodings are used. The encoder produces contextualized embeddings for every position, and we retain the $t$ target embeddings for the prediction.
@@ A.3 Architecture Experiments
For the fusion block, we consider two simpler alternatives to the attention mechanism, both of them underperforming: (1) Concatenation, by concatenating the semantic and numerical $d$-dimensional vectors into a $2d$-dimensional vector, and projecting them back via an MLP, and (2) Multiplication, by multiplying the semantic representation directly with the numerical value [^22] After rescaling it to be centered around 1 rather than 0, using a learned scaling factor. in a straightforward, parameter‐free manner as in [83, 86].
@@
Verbalization
We experiment with various verbalization strategies. For features, including the column name alongside the value consistently outperforms using the value alone. For target-aware tokens, we find that explicitly verbalizing all possible target values yields better performance than omitting them. In contrast, adding a dataset-level description prefix to improve contextualization offers no measurable advantage. Interestingly, TabSTAR maintains high performance even when no semantic metadata is present. We see this robustness as a strength.
@@ B.2 Finetuning
We finetune downstream tasks using LoRA’s implementation of the peft package [^24] https://github.com/huggingface/peft. We use a rank of $r=32$, set $\alpha=2r=64$ and $dropout=0.1$. We employ the same scheduler as in the pretraining phase, with the only difference being that we set $lr=0.001$, and increase the patience parameter for early stopping to 5. We apply a train-test split of 90%-10% and sample a validation set of 10%. As opposed to the pretraining, all batches are drawn from the same dataset. Therefore, we observe no effect from changing the mini-batch size when keeping the global batch size fixed to 128. We tune 1,597,440 out of TabSTAR’s 47,263,874 parameters (3.4%), spanning all blocks of the architecture. The frozen layers of the text encoder do not receive LoRA adapters.
@@ E.2 Baselines Implementation and Hyperparameters
CARTE natively supports textual inputs, and the TabPFN-v2 API client [^28] https://github.com/PriorLabs/tabpfn-client does as well, although its implementation details remain undisclosed. As all other baselines do not natively support free-text features, [^29] CatBoost includes a built-in text module, but it underperforms compared to dense text embeddings. we preprocess these features into fixed-size embeddings using skrub’s TextEncoder, which internally applies a frozen e5-small-v2 encoder to each semantic column. This aligns with the encoder used in TabSTAR, enabling a fair comparison across models. There are, however, two key differences in how TabSTAR handles these embeddings: First, the embeddings are specifically finetuned for the task, contributing significantly to its strong performance as shown in §5 and further analyzed in §G.2. The second detail is that skrub applies dimensionality reduction to 30 dimensions, as proposed by [32]. This compressed representation performs comparably to the full embedding space, while offering improved inference efficiency.
Hyperparameter Tuning
For the tuned models we use the Optuna package [^30] https://pypi.org/project/optuna/ with random search, with a budget of 4 hours for every run and parallelizing trials on 8 CPU cores. We use 5-fold cross-validation and take the best configuration selected based on this mean score. We use it then to retrain the model on the full training data.
@@ E.2.1 TabPFN-v2
E.2.2 TabICL
We run TabICL using its package [^32] https://pypi.org/project/tabicl/ with the default configuration. Since TabICL can only be evaluated on classification tasks, we exclude it from the regression analysis.
E.2.3 TabDPT
We run TabDPT using its package [^33] https://pypi.org/project/tabdpt/ with the default configuration. While it achieves state-of-the-art performance on some regression tasks, we encounter severe performance issues on others. Due to this inconsistency, we exclude TabDPT from the reported results and are in contact with the authors to investigate potential implementation issues.
E.2.4 CARTE